OpenStudy (anonymous):

A cat jumps straight up from the ground with an initial vertical velocity of 9 feet per second. a. Write an equation that models the height [h] (in feet) of the cat as a function of the time [t] (in seconds) since it jumps. b. After how many seconds does the cat land on the ground?

OpenStudy (saifoo.khan):

velocity = disp/time

OpenStudy (saifoo.khan):

velocity = h / t

OpenStudy (anonymous):

ho ho ho nice cat

OpenStudy (saifoo.khan):

LOL

OpenStudy (anonymous):

i think you are supposed to use \[h(t)=-16t^2+9t\]

OpenStudy (anonymous):

to find the amount of time the little kitty is in the air, set this equal to zero and solve. you get \[-16t^2+9t=0\] \[t(9-16t)=0\] \[t=\frac{9}{16}\]

OpenStudy (anonymous):

I see how you get t(9-16t)=0 But how does t=9/16?

OpenStudy (anonymous):

i set \[9-16t=0\] \[9=16t\] \[\frac{9}{16}=t\]

OpenStudy (anonymous):

What happens to the t in the beginning?